OpenStudy (anonymous):

What is the sum of an 8–term geometric series if the first term is 13 and the last term is –3,639,168?

OpenStudy (anonymous):

ok first term is \[a=13\] last term is \[13r^8=–3,639,168\] solve for r by dividing by 13 to get -279936 so \[r^7=-279936\] \[r=\sqrt[7]{-279936}=-6\] so now we have r

OpenStudy (anonymous):

i meant \[13r^7= –3,639,168\]

OpenStudy (anonymous):

so now use the formula for summing a geomtric series. i think it is \[\frac{a(r^8-1)}{r-1}\] but you should check it if so i get \[\frac{13((-6)^8-1)}{-7}\] lets see

OpenStudy (anonymous):

I got –3,119,285

OpenStudy (anonymous):

calculator give me \[-3119285\] so at least it is a nice integer

OpenStudy (anonymous):

me too!

OpenStudy (anonymous):

Yay! Thank you so much satellite :) you are my hero!! Could you help me with one more? What is the sum of a 6–term geometric series if the first term is 19 and the last term is 319,333?